LOG-IN
Displaying reports 1821-1840 of 3413.Go to page Start 88 89 90 91 92 93 94 95 96 End
R&D (FilterCavity)
Print this report.
EleonoraCapocasa - 11:55, Tuesday 26 November 2019 (1886)Get code to link to this report
Filter cavity remote lock

[Eleonora, Federico, Matteo]

The zero-detect board of the rampeatuo (Pic1) reads a signal (usually the cavity transmission) and compares it with a treshold signal in order to engage the lock. The threshold signal can be manually set with a potentiomenter and goes from -15 V to 15 V. Pierre installed a probe to monitor this threshold some times ago.

In the current configuration the rampeauto sums the transmission signal and the threshold and engages the lock if this sum is > 0. This means that if we don't connect the transmission signal, the cavity gets locked when we set the threshold above 0.  

By connecting the transmission and keeping the threshold below zero we can assure the the lock is engaged only when the transmission is higher than the absolute value of the treshold. This means that we can prevent the servo from locking on HOMs. In the current configuration, the cavity transmission (when it is locked and well aligned) is ~1.5 V. The threshold is set at -0.5 V so that the lock is engaged only when the transmission is higher than 0.5. 

In order to remotely control the lock, we used a stanford to subtract an offset (generated by the recenty installed DAC) to the transmission signal, before the rampeauto. If such offset is 0 the cavity stays locked, if it is larger than 1 V the trasmission signal sent to the rampeauto is lower that 0.5 V and the servo stops the lock. We set the offset at 1.5 V and we verified that the cavity lock can be controlled by adding and removing it.

I added a button on the main MEDM screen to control such offset (pic 2). From now on please try to use remote lock of the filter cavity as much as possible and keep the threshold knob where it is.

Images attached to this report
1886_20191126035450_zerodet.jpeg 1886_20191126035508_remote1.png
R&D (FilterCavity)
Print this report.
YuhangZhao - 00:06, Tuesday 26 November 2019 (1884)Get code to link to this report
The peaks remained in squeezing spectrum after the installation of new FI

The attached figure shows the remained peaks in the squeezing/anti-squeezing spectrum.

In the anti-squeezing spectrum, we could see that all the peaks appear at a harmonics of 50Hz.

In the squeezing spectrum, the peaks appear mainly at harmonics of 50Hz. Only 3 out of 14 peaks are not harmonics of 50Hz but they are close.

Images attached to this report
1884_20191125160603_peaksanti.png 1884_20191125160841_peakssqz.png
R&D (FilterCavity)
Print this report.
YuhangZhao - 22:10, Monday 25 November 2019 (1883)Get code to link to this report
Replacement of IR phase shifter

Eleonora and Yuhang

After Takahashi-san glued the PZT on the new mount. We soldered the PZT to a BNC connector and replaced it with the old IR phase shifter. (attached figure 1)

The alignment to IRMC was recovered. Then we locked it and tried to see the effect of scanning IR phase shifter. We sent a sine wave to IR phase shifter from 25-125V, which is almost the same output range of our servo. Then we found IRMC transmission is modulated by almost 12% in the pitch direction. (attached figure 2)

The reason why pitched is mainly modulated is guessed as that the mirror on top of PZT is a bit heavy so it has some pitch tilt, the PZT motion creates mainly pitch misalignment. Also, the beam is tilted hitting on IR phase shifter is tilted in the pitch direction.

Images attached to this report
1883_20191125140724_wechatimg598.jpeg 1883_20191125140731_wechatimg599.jpeg
R&D (FilterCavity)
Print this report.
NaokiAritomi - 19:09, Monday 25 November 2019 (1882)Get code to link to this report
FC IR TRA channel is not working well

FC IR TRA channel (ADC CH1) is not working well. We used ADC CH16 (FDS-SEISM_Z_IN1) for FC IR TRA instead.

KAGRA MIR (Absorption)
Print this report.
SimonZeidler - 18:53, Monday 25 November 2019 (1881)Get code to link to this report
PCI: Measuring the beam-profile of the laser after the telescope

Pengbo, Simon

Today, with the very much appreciated help from Manuel, we could start the translation-scans of the razor blade in horizontal direction at different Z-values in order to analyze the beam-profile automatically.
For the measurements itself, the powermeter is connected via a BNC cable to the DC-port of the LabView program.

Pengbo has written a Python-script to fit several of those scans at once so that we can easily get the results.

The laser-power of the pump beam is set to be ~110 mW for the entire measurements.

We started with our measurements actually in the area where the waist of the beam is supposed to be (Z-distance = 2mm, dZ = 0.1 mm, Zcenter = 74.1).
The next set of measurements will be over a wider range in Z further away from the waist.

R&D (FilterCavity)
Print this report.
RyutaroTakahashi - 18:04, Monday 25 November 2019 (1880)Get code to link to this report
Comment to Fail of TMP in south end (Click here to view original report: 1878)

Baking time today (11/25): 14:00-17:00 (3 hours)

R&D (Cryogenic)
Print this report.
SatoshiTanioka - 17:23, Monday 25 November 2019 (1879)Get code to link to this report
The vacuum gauge's problem

[Tomaru-san, Tanioka]

Today, we investigated the problem of the vacuum gauge.
Tomaru-san cleaned the inside of it, then I installed it to the chamber.
We checked the behavior by pumping down, but the situation was not improved.
Around 1.2*10-2 Pa, the vacuum gauge showed an error related to the sensor.

We will send it to the company and ask for repairing or buy a new one.

Comments related to this report
SatoshiTanioka - 15:58, Tuesday 26 November 2019 (1889)

I removed the gauge from the chamber.

R&D (FilterCavity)
Print this report.
RyutaroTakahashi - 15:39, Monday 25 November 2019 (1878)Get code to link to this report
Fail of TMP in south end

I found that the TMP in south end was failed. The TMP controller showed red indication but the LCD was disappeared.

Similar to the case in BS, oil vapor went from the RP to the TMP. Therefore the RP was replaced to the dry pump from the MC.

The TMP controller was replaced to the other one from the BS. The TMP is under baking to remove the contamination due to the oil vapor.

Comments related to this report
RyutaroTakahashi - 18:04, Monday 25 November 2019 (1880)

Baking time today (11/25): 14:00-17:00 (3 hours)

RyutaroTakahashi - 14:02, Tuesday 26 November 2019 (1887)

Baking time today (11/26): 9:30-10:30 (1 hour)

After the temparature settled, I opened the gate valves for the TMP. The pressure was decreased from 1x10-6 mbar to 2x10-7 mbar.

R&D (FilterCavity)
Print this report.
EleonoraCapocasa - 12:08, Monday 25 November 2019 (1877)Get code to link to this report
Rough actuators calibration for cavity mirrors

I tried to compute a rough calibration for the cavity mirrors actuation in DC.

I injected an offest of 10.000 counts in pitch and yaw, in INPUT and END mirrors, one at a time, and record the change seen by OPLEV.

Then I used the calibration reported in entry #1874 to convert it in urad:

10.000 counts injected 

INPUT

END

PITCH

63 urad

52   urad

YAW 220 urad 268    urad

Pitch is stiffer than yaw. There is a difference of 15/20% beetween INPUT and END that can come from both errors in OPLEV calibration and actuation unbalance. Also It is a bit strange to me that for the same amout of injected counts pitch motion is larger for INPUT and yaw motion is larger for END.

R&D (FilterCavity)
Print this report.
EleonoraCapocasa - 13:42, Saturday 23 November 2019 (1876)Get code to link to this report
DAC troubles solved

[Federico, Matteo, Eleonora]

In the past few days we had many troubles with our DACs. Several channels were not working and moreover the issue seemed somehow "non stationary": some channels stopped working, than worked again, etc..

After we veryfied that the connections and the hardware have no evident problems, we decided to change the old DAC board with a new one. This seems to have fixed the problem. We also installed the cable to connect the second DAC board to the timing box and test the signal chain. We had to deal with several issues:

1) Dsub to BNC board installed in the cleanroom had a strange behaviour. We found some jumpers were missing inside and we soldered them (we could't find proper jumpers in the elec-shop). Now it is OK.

2) The AI board has only 8 channels (instead on 16)  but only channels 2-3-4 seem to work fine. Channel 1 has a large offset and the signal is very much attenuated, channels 5-6-7-8 show crazy things.

3) In general both the dsub to BNC boards show an oscillation at about 1 MHz, with amplitude of volts, when the output is probed with a long cable. This issue was also observed in KAGRA. (Entry #6266 of Klog). According to our expert electronician this is due to the cable impedence which induces an auto-oscillation of the last op-amp. This could have been solved by inserting a resistence before the output. Note that we are currently sending such dirty signals to our coil-drivers. Hopefully the oscillation frequency is so high that they don't really care? Note that most but not all the channels show this behaviour. 

R&D (FilterCavity)
Print this report.
YuhangZhao - 00:54, Saturday 23 November 2019 (1875)Get code to link to this report
QPD response to different size and power of 532nm light

Federico, Yaochin, and Yuhang

1. Test of small size beam with different power (attached figure 1 and 2)

power varies from 5mW to 12mW.

2. Test of 12mW light with a different beam size (attached figure 3 and 4)

3. The test of the large beam with different power (attached figure 5 and 6)

varies from 0.5mW to 12mW.

Note: According to the AA telescope design, the beam size is ~260um. The available power we can send to QPD is 2mW, the power distributed to rach quarter is 0.5mW.

Images attached to this report
1875_20191122165259_small1.png 1875_20191122165310_small2.png 1875_20191122165338_size1.png 1875_20191122165347_size3.png 1875_20191122165356_large1.png 1875_20191122165403_large2.png
R&D (FilterCavity)
Print this report.
EleonoraCapocasa - 19:36, Friday 22 November 2019 (1874)Get code to link to this report
OPLEV signal calibration

I summarize the calibration for the OPLEVs.

For the conversion from counts to Volts, I used the following value:  1 V = 1638 counts  ->  cal =  6.105e-4 [V/count].

Note that for BS and PR we use TAMA PSD and no lens for shift/tilt decoupling. 

 

PR (entry #276)  --- SUM:  11 V  

Yaw:   0.24+/-0.03 [mrad/V]  = 0.146+/-0.02 [urad/count]

Pitch:  0.33+/-0.03 [mrad/V]  = 0.21+/-0.02 [urad/count]

NB: for geometrical reasons pitch is multiplied by a factor sqrt(2) (see attachment in entry 276)

 

BS (entry #337)  --- SUM: 13.4 V

Yaw:   0.37+/-0.03 [mrad/V]  = 0.25+/-0.02 [urad/count]

Pitch:  0.37+/-0.03 [mrad/V]  = 0.25+/-0.02 [urad/count]

NB: Since in this case the incident angle is almost normal the factor sqrt(2) can be neglected.

 

INPUT (entry #278 and pag 168 of my Phd thesis)  --- SUM: 3.2 V

Yaw:   0.038 [mrad/V]  = 0.027 [urad/count]

Pitch:  0.062 [mrad/V]  = 0.038 [urad/count]

NB: We considered a factor 100 coming from SR560 amplifiers 

 

END (entry #278 and pag 168 of my Phd thesis)  --- SUM: 4.45 V

Yaw:   0.030 [mrad/V]  = 0.018 [urad/count]

Pitch:  0.043 [mrad/V]  = 0.026 [urad/count]

NB: We considered a factor 100 coming from SR560 amplifiers 

 

It would be good to implement real time calibration for OPLEV channels.

R&D (Cryogenic)
Print this report.
SatoshiTanioka - 18:58, Friday 22 November 2019 (1870)Get code to link to this report
Beam Jitter Suppression

Today I checked the beam jitter induced by the AOM frequency shift.
I measured beam shift by a beam profiler at ~2.5m distance from the AOM.
The beam shift was about 0.16mm by ~15MHz frequency shift and this is small enough for this experiment.

[note]
Tweaking the alignment of double-pass AOMs is needed when the beams are locked to the folded cavity in order to maximize the diffracted beam power.

R&D (FilterCavity)
Print this report.
RyutaroTakahashi - 18:06, Friday 22 November 2019 (1873)Get code to link to this report
Replacement of TMP

The TMP in BS was replaced to STP-1003 on the 21st. After the one-day monitoring, I opened the gate valve on the pump. The pressure in the arm input was improved from 1.7x10-7 mbar to 8.5x10-8 mbar. Unfortunately the vacuum gauge on the BS is not working due to the over life-time.

R&D (FilterCavity)
Print this report.
CamillaDerossi, Fiori, Paoletti - 15:56, Friday 22 November 2019 (1868)Get code to link to this report
tapping test on SQZ bench
We installed our Trillium seismometer on the floor (next to the TAEC, picture 1) and the Meggit accelerometers on the SQZ bench along the south edge (picture 2). Data were acquired with the Centaurus.
We performed some tapping test in order to identify the resonances of the bench. On the attached plots x stands for the NS direction, y for the EW and z for the vertical.

- when tapping along NS we excited the 10,4 Hz peak (plot 3 shows the spectra during the tapping along the 3 directions)
- when tapping along EW we excited the 14,2 Hz peak (plot 4)
- when tapping along the vertical direction we excited a broad peak at 41.8 Hz and another broad bump around 84Hz, but not very clear (plot 5)

We also performed a switch off test of the fans over the SQZ bench. As you can see from plot 6, it doesn't seem to make any difference along the z direction

The last plot shows the FFTs of the accelerometers on the SQZ bench while excited and the ones of the seismometer on the floor. All 3 directions are shown together. It would be useful to compute the transfer functions between them.

In conclusion: 10,4 and 14,2 Hz are resonances of the SQZ bench, and the fan over it doesn't seem to create any excitation.
Images attached to this report
1868_20191122055624_picturetrilliuandteac.jpeg 1868_20191122055738_pictureaccbench2.jpeg 1868_20191122060645_benchtappingns.png 1868_20191122060659_benchtappingew.png 1868_20191122060715_benchtappingvert.png 1868_20191122061547_switchofffansqzbench.png 1868_20191122061613_fftsbenchtappingvsground.png
R&D (Cryogenic)
Print this report.
SatoshiTanioka - 12:01, Friday 22 November 2019 (1869)Get code to link to this report
Comment to The behavior of vacuum gauge (Click here to view original report: 1863)

I checked the output voltage from the vacuum gauge and it was 0.2V.
This indicates that there is an error in the B-A gauge or the sensor.

I did degas, but this not change the situation.

One possible solution is to clean up the gauge.

R&D (FilterCavity)
Print this report.
YuhangZhao - 11:28, Friday 22 November 2019 (1867)Get code to link to this report
Matching between OPO and AMC

Yaochin and Yuhang

The spectrum of BAB into AMC is as following. TEM00: 5.56V, Higher-order 1: 6.8mV, Higher order 2: 10.6mV.

So the matching is 5.56/(5.56+6.8e-3+10.6e-3) = 99.69%

Images attached to this report
1867_20191122032531_tek00075.png 1867_20191122032542_tek00076.png 1867_20191122032550_tek00077.png
R&D (FilterCavity)
Print this report.
NaokiAritomi - 00:35, Friday 22 November 2019 (1866)Get code to link to this report
Comment to The characterization of squeezing after the optimization of homodyne (Click here to view original report: 1865)

Did you tune CC2 demodulation phase for each green power?

R&D (FilterCavity)
Print this report.
YuhangZhao - 23:57, Thursday 21 November 2019 (1865)Get code to link to this report
The characterization of squeezing after the optimization of homodyne

Yao-Chin and Yuhang

For the optimization of homodyne, we basically align again the homodyne. Including the matching of IRMC/AMC and OPO/AMC, the balance of homodyne, the beam height regulation.

The situation of IRMC/AMC matching is 99.95% while OPO/AMC is 99.7%.

The balance of homodyne now is done without aligning its BS but by aligning the lens before homodyne.

Benefit from not aligning homodyne BS, we could make squeezing go to homodyne flatly. Also, in this case, the balance of homodyne is easier.

After this work, we characterize squeezing again. We measured squeezing and anti-squeezing for power from 20mW to 60mW with an interval of 5mW. The result is shown in the attached first figure. 

In the attached figure 2, we see there was a peak appearing. I need to mention that this peak appears after we did several measurements.

By fitting the squeezing and anti-squeezing, we could find out the loss and phase noise information. However, this time, the fitting and raw data has a larger discrepancy.

#############

I am sorry that I forgot to put the information about the Green power and demodulation phase.

green pump power MZ offset OPO temperature p-pol locking frequency CC2 demodulation phase(sqz) CC2 demodulation phase(asqz)
20 4.1 7.166 175 75 160
25 4.2 7.167 175 90 170
30 4.3 7.18 175 100 175
35 4.4 7.18 175 120 170
40 4.5 7.19 180 120 160
45 4.6 7.19 170 125 155
50 4.7 7.19 150 125 160
55 4.8 7.195 150 125 155
60 4.9 7.2 150 135 155

#############

In the last attached figure, we can see the fit without considering the sqz-asqz for 55, 60mW.

Note: the anti-squeezing level is lower than the measurement did last week. This may come from the alignment issue of green pump.

Images attached to this report
1865_20191121154504_figure1.png 1865_20191121155736_peak.png 1865_20191121155743_fit.png 1865_20191122050938_fit2.png
Comments related to this report
NaokiAritomi - 00:35, Friday 22 November 2019 (1866)

Did you tune CC2 demodulation phase for each green power?

R&D (FilterCavity)
Print this report.
YuhangZhao - 23:34, Thursday 21 November 2019 (1864)Get code to link to this report
Two lens before homodyne are scratched

Yaochin and Yuhang

We found the two lenses before homodyne both have a scratch. One has scratch just in the center area. Another has a larger area scratched but not in the center. 

The photo of the lens is shown in the attachment.

We will order new lenses.

Images attached to this report
1864_20191121153352_wechatimg596.jpeg 1864_20191121153400_wechatimg597.jpeg